OpenStudy (anonymous):

Find the area under the curve y = 1/(3 x^3) from x = 1 to x = t and evaluate it for t = 10, t = 100. Then find the total area under this curve for x ≥ 1. a) t=10 b)t=100 c) Total area

OpenStudy (anonymous):

Area under a curve \[\int\limits_{1}^{t}1/(3x^3)\] Set t= 10 then solve.\[\int\limits_{1}^{10}1/(3x^3)\] Set t=100 then solve.\[\int\limits_{1}^{100}1/(3x^3)\] They are looking for the absolute value of the integrand, if I'm not mistaken, for total area. So it would look like. Then solve and repeat. \[\int\limits_{1}^{10}|1/(3x^3)|\]

OpenStudy (anonymous):

Let me know if you want me to show you how to take the integral by hand.